OpenStudy (anonymous):

Solve the equation by completing the square. 4x2 − 8x − 1 = 0 What is: x = (smaller value) x = (larger value)

OpenStudy (anonymous):

x=1+(5/4)^1/2 x=1-(5/4)^1/2

OpenStudy (anonymous):

how do you treat that exponent

OpenStudy (anonymous):

is the answer - 2 and 2